EPS-HEP99 Abstract # 000 Parallel sessions: Plenary sessions:

Measurement of Differential Jet Production in Deep-Inelastic Scattering at High Q^{2} at HERA

H1 Collaboration,

DESY, Notkestr.85, D-22607 Hamburg, Germany

Jet production in neutral–current DIS at HERA is studied at high values of 150 < Q^{2} < 5, 000 GeV^{2} using data collected with the H1 detector in 1995-1997. Measurements of jet angular distributions, differential (1+1)- and (2+1)-jet rates, jet–jet invariant masses up to 200 GeV and other jet variables are presented. The large number of events allows us to study the Q^{2} dependence of the measured distributions. The mentioned distributions are determined for several alternative jet algorithms, and are compared with perturbative QCD predictions in NLO and with QCD Monte Carlo models. The value of the strong coupling constant α_{s} is determined.